### v1

# Out of Sight, Still in Mind: Reasoning and Planning about Unobserved Objects with Video Tracking Enabled Memory Models

Abstract
Robots need to have a memory of previously observed, but currently occluded objects to work reliably in realistic environments. We investigate the problem of encoding object-oriented memory into a multi-object manipulation reasoning and planning framework. We propose DOOM and LOOM, which leverage transformer relational dynamics to encode the history of trajectories given partial-view point clouds and an object discovery and tracking engine. Our approaches can perform multiple challenging tasks including reasoning with occluded objects, novel objects appearance, and object reappearance. Throughout our extensive simulation and real-world experiments, we find that our approaches perform well in terms of different numbers of objects and different numbers of distractor actions. Furthermore, we show our approaches outperform an implicit memory baseline.
I Introduction
For robots to assist humans in their daily lives in roles such as home assistants and caregivers for elders, they must be able to reason about objects not observed in their current perceptual data. For example, if asked to retrieve an apple stored inside a cabinet, the robot should remember where it was placed and that it must first open the cabinet to grasp the object. Further if the robot picks up a box containing objects it previously placed inside the container, it should know that these objects will move with the box, while those sitting near the box will not. We desire for robots to maintain these models over many separate task assignments from their operators in order to enable long-term autonomy.
The problem of maintaining a persistent model of a robot’s world from partial observations has a rich history in robotics, primarily in the form of mapping [1, 2, 3, 4, 5]. While some attempts have been made to incorporate object-level semantics into sensor-derived maps [6, 7, 8] these approaches primarily build immutable, monolithic models not suitable for complex manipulation planning such as multi-object rearrangement [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Such reasoning can enable robots to predict preconditions for actions [12], achieve language goals [13], and achieve logical goal relations [18, 15]. Thus a problem arises in how to build a consistent estimate of the world state amenable to contemporary, learning-based approaches to manipulation planning.
In contrast, the standard learning-based approach to persistent memory treats the problem as a sequence prediction task, typically using autoregressive recurrent or transformer-based neural networks [19, 20, 21, 22, 23]. While these models have shown some interesting results, we hypothesize that they will have difficulty maintaining long-term memory of objects that remain occluded through many updates to a recurrent memory. One approach to overcome this issue would be to maintain the entire history of robot observations in order to use as input to a model (e.g. transformer) that could selectively attend to the relevant bits of input to determine a coherent state estimate. This is obviously untenable for long-horizon tasks as the input would grow linearly with time, inevitably becoming too large to efficiently manage.
Figure 2: Overview of our approaches. As the robot takes action over time, some objects may disappear and reappear, and some objects may newly appear in the scene. In this paper, we propose two types of object-oriented memory, called DOOM and LOOM, that enable the robot to plan with occluded and newly appeared objects. DOOM and LOOM utilize a UVOS algorithm to keep track of the current object list and update object memory slots based on the occlusion status of each object accordingly. Best viewed in color.
We instead argue for an explicit way to encode the state of all currently and previously observed objects in a consistent manner. In recent years, there have been promising solutions for the unsupervised video object segmentation (UVOS)[24] task where the algorithm simultaneously discovers previously unknown objects and tracks them through time [25, 26, 27, 28]. Especially interesting for our sake, such algorithms allow for tracking under heavy and long-term complete occlusions. This opens up an opportunity to use a UVOS algorithm in robot manipulation tasks where we may not have segmentation labels for objects of interest.
In this paper we advocate for using a UVOS algorithm to explicitly manage our object-oriented memory. We hypothesize that explicit encoding of objects into the memory will be more robust than implicit, autoregressive models in managing long-term history and successfully performing downstream planning. We desire that the representation provide sufficient information to enable prediction of inter-object and object-environment relations, while also having the ability to predict the effects of the robot’s actions on all objects in memory.
Specifically, we examine incorporating an explicit UVOS-based memory model into the framework from Huang et al. [17, 18], which was an effective framework to learn relational dynamics across varying object and environments. Key to its success is the ability to encode a variable number of objects for a given observation using a graph neural net [17] or transformer-based encoder [18]. Prediction of relations enables the model’s use in logic-based task planning [29], where relations have proved an effective means of communication between robots and humans [15, 13, 10, 30, 12, 17]. However, the existing framework assumes all relevant objects to be observable; and thus fails to successfully plan and execute in the scenarios shown in Fig. 1.
We propose two ways to integrate predictions from the video tracker of [28] with the relational dynamics prediction model of [18]. Both approaches augment the current state estimate with information from currently unobserved objects for use in predicting inter-object relations and action effects. One approach directly augments the latent space of the dynamics model by concatenating the previously predicted latent state tokens for unobserved objects with those currently observed. We term this Latent Occluded Object Memory (LOOM). The second method, termed Direct Occluded Object Memory (DOOM), directly augments the input point cloud with the previously observed object point cloud transformed based on its previously predicted pose estimate. Figure 2 illustrates the LOOM and DOOM models.
To validate our proposed approaches, we test on varying tasks involving occluded objects, novel objects, and objects that reappear. We show some example tasks in Fig. 1. We show that both forms of memory-based state augmentation can reliably plan actions to change inter-object relations involving currently unobserved objects, such as placing object $A$ to the left of currently occluded object $B$. Furthermore we compare our explicit memory approach to a baseline implicit memory model in a scenario specifically designed to test the models’ long-term memory. Our experiments show that the implicit memory model fails to effectively reason about occluded objects, while our explicit, UVOS-based memory models continue to plan reliably. Our work thus represents the first successful approach of using memory models for reasoning about and planning to inter-object goal relations for an a priori unknown, variable number of objects. For more real-world robot executions, additional results, and supplemental material, see our website https://sites.google.com/view/rdmemory.
II Related work
Reasoning about object permanence is an important capability for robot manipulation [31, 22, 32, 33]. Xu et al. [31] and Ebert et al. [22] are the first ones to propose deep learning models to reason about occluded objects and do downstream planning with the learned dynamics models. However, both of these approaches assume goal images for planning which may not always be available from human operators. They additionally examine only planar pushing tasks with only a single moving object at each time. Our work examines a much more diverse set of tasks and skills, while also requiring only logic-based goal representations. Curtis et al. [33] propose a system with modules that can estimate affordances and properties to perform multi-step manipulation tasks with unknown and occluded objects. However, they assume complete object shape and have many engineered modules including the affordance module.
Encoding memory for manipulation and navigation has received some attention [34, 23, 35, 36, 37]. Shafiullah et al. [34] propose an implicit model for semantic navigation but this model cannot be directly applicable to manipulation tasks. Kim et al. [23] propose both RNNs and transformers to predict memory-based gaze via imitation learning. However, the work focuses on gaze prediction and shows no experiments with occluded objects or objects with novel appearances.
Reasoning about occluded objects with memory has received attention in the vision community in the context of visual tracking [38, 39, 28] and video object segmentation [40, 41]. However, these memory models have been studied in the context of object re-identification across occlusion only (i.e., appearance matching) and cannot be utilized for reasoning about how occluded objects may move based on a robot’s actions. Our work builds on these approaches by combining them with a learned action effects model.
Figure 3: Two examples of our training dataset (top row) and one testing example (bottom row). We train with a maximum of 5 segments including objects and the environments. The testing example has 8 segments with different shapes and a novel view point. In the history, the robot pushes the mug below the shelf then picks and places two apples inside the bowl. During planning, the robot picks and places the orange to achieve the goal relation based on the current observation and history. Left/Right are defined from the robot’s viewpoint.
Other works have also investigated using graph neural networks [42, 17, 10, 43, 44], transformers [13, 45, 46, 47, 48], and diffusion models [9, 49, 50] to reason about and manipulate multiple objects. However, none have examined explicit memory management. Since multi-object manipulation usually requires a multi-step process, task and motion planning has shown promise in solving long-horizon, multi-object planning problem [51, 52, 53, 54, 18, 55]. These planners either require all objects to be known a priori in order to perform tracking for belief space planning [53] or ignore occlusion of unknown objects.
III Approach
We assume the robot perceives the world as a point cloud, $Z_{t}$, at each timestep $t$. The robot then takes action $A_{t}$ and receives subsequent observation $Z_{t+1}$. At new observation $Z_{t+1}$, some objects may have become occluded and other, new objects may appear. Based on the history of observations and actions $Z_{0:t}$, $A_{0:t}$ we would like the robot to plan to achieve a goal, potentially involving previously observed, but currently occluded objects.
We define the goal as a logical conjunction of $M$ desired object and environment relations, $\mathbf{g}=r_{1}\land r_{2}\land...\land r_{M},r_{j}\in\mathcal{R}$, where $\mathbf{g}$ denotes the goal conjunction, $r_{j}$ represents each goal relation, and $\mathcal{R}$ denotes all possible relations. Our robot is given a set of $L$ parametric action primitives $\mathcal{A}=\{A_{1},\ldots,A_{L}\}$ where $A_{l}$ defines a skill, which has associated continuous skill parameters $\theta_{l}$. For example, a push skill is defined with parameters encoding the push direction and length, or a pick-and-dump skill is defined with parameters encoding the grasp pose and dump pose. The robot’s planning task is defined as finding skills and skill parameters $\tau=((A_{0},\theta_{0}),\ldots,(A_{H-1},\theta_{H-1}))$ that, when sequentially executed, transform the objects so they satisfy all desired object and environment relations in the goal $\mathbf{g}$.
To solve this problem we propose a novel memory-based neural network framework. Instead of taking the entire history of observations as input, the model takes the current observation, $Z_{t}$, current action $(A_{t},\theta_{t})$ and a compressed memory of the previous observations $Z_{0:t-1}$ and actions $A_{0:t-1}$, and predicts the resulting relations $r_{t+1}^{\prime}$ and object poses $p_{t+1}^{\prime}$. As shown in Fig. 1, this enables our framework to remember the pose of disappeared objects after several actions. By chaining together predictions, we can effectively perform multi-step planning. We propose two different implementations of this framework called DOOM and LOOM, which respectively use a point cloud-based encoding and latent space encoding to represent the memory, $Q$. We show an overview of our approaches in Fig. 2. We now explain the various components of this framework.
Segmentation and Tracking We first perform UVOS using [28] and obtain $N$ object and environment segments $O^{i}_{t}\subset Z_{t},i=1,2,...,N$ for a timestep $t$. The UVOS approach checks for new objects by examining the objectness of object proposals on key frames, when the robot finishes the execution of a skill. Then object segments that can be matched with previously tracked objects are assigned consistent IDs, while those that do not match any previous object are assigned new IDs. For non-key frames, the algorithm tracks all segments with a space-time transformer model [41] where each object has its own memory. It stores prior appearances of all objects hence can easily re-identify previously observed and disappeared objects. To familiarize the UVOS model with the robot, it is fine-tuned using annotations of the robot arm and objects from YCB-Video dataset [56]. Nonetheless, it still demonstrates generalization to novel object types, like the tall cups (Row. 1), apple and orange (Row. 2) shown in Fig. 1 that it was never trained on. For environment segments, we use RANSAC [57] to find planar surfaces.
Pose and relation detection To jointly reason about the pose of each object and the relations between each pair of objects, we first process each segment $O^{i}_{t}$ with a point cloud encoder [58] to get a feature vector per segment with $P^{i}_{t}=E_{p}(O^{i}_{t})$. We use a learned positional embedding to encode the ID of each object as $I_{i}$. We randomly generate the ID for each object during training to improve generalization [59]. We get the latent space at current step $t$ as $\mathbf{x}^{i}_{t}=E(P^{i}_{t}\oplus I_{i})$, where $E$ is a transformer encoding the interaction between different segments. To predict object pose and relations at the current step, we define three different decoders: the pose decoder $D_{p}$, the relation decoder $D_{r}$, and the environment identity decoder $D_{e}$. The third decoder classifies whether a given segment is movable or not (e.g. is a table or shelf) as $\hat{e}^{i}_{t}=D_{e}(\mathbf{x}^{i}_{t})$. The associated predictions of the other decoders are $\hat{p}^{i}_{t}=D_{p}(\mathbf{x}^{i}_{t})$ and $\hat{r}^{ij}_{t}=D_{r}(\mathbf{x}^{i}_{t},\mathbf{x}^{j}_{t})$.
TABLE I: Comparison in terms of the F1 score on the relational predictions.
Objects
4
5
6
7
all
Distractors
1
2
3
all
DOOM
0.998
0.975
0.974 0.958 0.976
DOOM
0.918 0.907 0.901 0.909
LOOM
0.994
0.978
0.972
0.951
0.974
LOOM
0.894
0.873
0.868
0.878
Baseline
0.938
0.786
0.765
0.702
0.798
Baseline
0.797
0.763
0.751
0.770
Pose and relation prediction To get the predicted pose and relations for future steps, we utilize a latent space dynamics model to propagate the current state information to the future after multiple actions. We define the latent space dynamics to propagate the $\mathbf{x}^{i}_{t}$ to ${\mathbf{x}^{\prime}}^{i}_{t+1}$ as ${\mathbf{x}^{\prime}}^{i}_{t+1}=\delta(\mathbf{x}^{i}_{t},A_{t}(i,\theta))$. where $i$ is the ID of the object to manipulate and $\theta$ are continuous parameters parameters. We use the same learned positional embedding to encode the discrete ID, as well as an MLP $M_{\text{ac}}$ to encode $\theta$. Since we have different skills, we use a different $\delta_{l}$ and $M_{\text{ac}_{l}}$ for each skill parameter $l$. After we get ${\mathbf{x}^{\prime}}^{i}_{t+1}$, our framework decodes it to a predicted pose ${p^{\prime}}^{i}_{t+1}$ and a relation ${r^{\prime}}^{ij}_{t+1}$. Note we use ${p^{\prime}}^{i}_{t+1}$ which discriminates from $\hat{p}^{i}_{t+1}$, as ${p^{\prime}}^{i}_{t+1}$ comes from the observation $Z_{t}$ while $\hat{p}^{i}_{t+1}$ comes from the observation $Z_{t+1}$. With the ability to predict the pose and relation with latent space dynamics, even if there are occluded objects in $Z_{t+1}$, we can still estimate the pose of the occluded objects from $(Z_{t},A_{t},\delta)$. Finally, we can apply $\delta(\cdot)$ $H$ times with a sequence of actions ${A_{0},...,A_{H-1}}$ to predict states $H$ time steps ahead as ${p^{\prime}}^{i}_{t+H}$ and ${r^{\prime}}^{{ij}}_{t+H}$.
Reasoning about occluded objects We explore two different approaches to reason about occluded objects. Consider object $k$ to be occluded. In our first approach, we use the predicted pose to transform the point cloud of object $k$ from time $t-1$, $O^{k}_{t-1}$, to $t$ in order to recover the missing observation as ${O^{\prime}}^{k}_{t}$. We then combine the transformed point cloud ${O^{\prime}}^{k}_{t}$ with the current observations as $Z^{\prime}_{t}={O^{1}_{t},...,{O^{\prime}}^{k}_{t},...,O^{N}_{t}}$. This allows the relational classifier to detect the relations at time $t$ even if the $k$th object disappears from the scene. We can repeat this process for any arbitrary number of occluded objects. We name this DOOM, for Direct Occluded Object Memory. We can alternatively copy the predicted latent space embedding ${\mathbf{x}^{\prime}}^{k}_{t}$ for the occluded object to the current latent state as $\mathbf{x}^{k}_{t}$ giving updated latent state $\hat{\mathbf{x}}^{\prime}_{t}=\mathbf{x}^{0}_{t},...{\mathbf{x}^{\prime}}^{k}_{t},...,\mathbf{x}^{N}_{t}$. We call this approach LOOM for Latent Occluded Object Memory. Figure 2 illustrates DOOM and LOOM.
Reasoning about reappeared objects When objects become hidden, we either augment the point cloud space or the latent space with DOOM or LOOM. When occluded objects reappear, for example, the $k$th object reappears at a timestep $t$, the UVOS tracker identifies this as the same ID for the previously occluded object. We can then remove the augmented memory state associated with object $k$ from the model and instead pass this object’s observation through the point cloud encoder like normal for both DOOM and LOOM.
Reasoning about novel objects’ appearance UVOS identifies novel objects when they first appear. In this case, each novel object segment receives a unique ID and is handled like any other object currently observed in the point cloud.
Planning with DOOM and LOOM For planning we use the cross-entropy method (CEM) as in [17, 18]. Since planning takes place prior to future observations, nothing changes when using DOOM or LOOM compared to the model in [18], except that the initial state must encode any unobserved objects as described above.
Training details Our training loss is the sum of three terms. First, we want our model to predict the pose, relations, and environment identity at the current step $t$. Given prediction $(\hat{p}^{i}_{t},\hat{r}^{ij}_{t},\hat{e}^{i}_{t})$ and ground truth labels $(p^{i}_{t},r^{ij}_{t},e^{i}_{t})$, we get the current observation loss $L_{c}=\sum_{t=1}^{H}CE(\hat{r}^{ij}_{t},r^{ij}_{t})+||\hat{p}^{i}_{t}-p^{i}_{t}||_{2}^{2}+CE(\hat{e}^{i}_{t},e^{i}_{t})$. Second, we require our models to learn that the predicted latent space from the previous observation ${\mathbf{x}^{\prime}}^{i}_{t+x}$ should be similar to the latent space from the current time observation $\mathbf{x}^{i}_{t+x}$. We call this the latent space regularization loss, defined as $L_{ls}=\sum_{t=1}^{H}\sum_{x=t}^{H}||\mathbf{x}^{i}_{t+x}-{\mathbf{x}^{\prime}}^{i}_{t+x}||_{2}^{2}$. Last, our model should predict correct outputs from the state, ${\mathbf{x}^{\prime}}^{i}_{t+x}$ predicted by the latent dynamics; giving loss: $L_{d}=\sum_{t=1}^{H}\sum_{x=t}^{H}CE({{r^{\prime}}^{ij}_{t+x}},r^{ij}_{t+x})+||{p^{\prime}}^{i}_{t+x}-p^{i}_{t+x}||_{2}^{2}+CE({e^{\prime}}^{i}_{t+x},e^{i}_{t+x})$. We train on the full loss function $L=L_{c}+L_{ls}+L_{d}$ using the Adam optimizer.
IV Experiments & Results
Figure 4: Execution success rate model comparison as a function of (left) the number of objects in the scene, and (right) the number of distractor actions. We find that our approaches perform consistently well across both conditions, outperforming the baseline by a large margin. This is especially prominent in terms of the numbers of distractor actions. The legend applies to both plots.
TABLE II: Real World Planning Success.
Objects
4
5
6
7
all
Distractors
1
2
3
all
DOOM
5/5
5/5
5/5
4/5
19/20
DOOM
5/5
4/5
5/5
14/15
LOOM
5/5
4/5
5/5
5/5
19/20
LOOM
4/5
5/5
4/5
13/15
[17, 18]
0/20
[17, 18]
0/15
IV-A Dataset Generation and Relations Definition
We generate a dataset with more than 20,000 skill executions. We generate skill executions with three steps: (1): We create scenes with a variable number of objects in the Isaac Gym simulator [60]. (2): We generate a random skill primitive with random continuous parameters. (3): We execute this skill in simulation. We repeat this process to create a training sequence with multiple timesteps. After the skill executions, we record the point clouds, poses, bounding boxes, and relations pre and post-manipulation. During data collection, if an object $O^{k}_{t}$ disappears at a timestep $t$ due to occlusion, we augment the point clouds at $t$ with the transformed point cloud of the disappeared object from a previous timestep. We show an example of the dataset generation process in Fig. 3. In this section, the number of objects is defined as the number of segments including objects and environments.
We define 9 relations for model evaluation: left, right, front, behind, above, below, contact, boundary, inside. The relations are defined based on the pose and extent of object bounding boxes in the simulation. The simulator provides us with the contact relation. We define inside(A, B) = 1 if the bounding box of A is surrounded by the bounding box of B, except that the inside object can stick out the top of the larger object a maximum of 3cm. We define the boundary relation as in [18] and other relations (left, right, etc.) as in [15]. Furthermore, we use the inside relation as a heuristic during planning and don’t attempt to directly grasp or push objects predicted to be contained inside others.
IV-B Baseline approach
We use a transformer to implicitly encode the history similar to [19] as our baseline approach. The transformer learns how to combine the history of previous observations and the current observation. For each observation $Z_{t}$, we concatenate the accumulated history with a type embedding $h_{type}$ and feed it into an MLP to get $h_{t}=MLP([E_{p}(E(X_{t}))+h_{t-1},h_{type}(t)])$ while we use $h_{0}=E_{p}(E(X_{0}))$. The type embedding $h_{type}$ encodes whether the segment is observable at the current step. We pass $h_{t}$ to the transformer to model the interaction between object tokens, and then use the decoders to predict object relations and poses.
IV-C Relational prediction evaluation
To test our hypothesis that an explicit object-oriented memory representation is more effective, we evaluate the ability of DOOM and LOOM to predict relations after an action. We define a distractor action as an action applied after some objects disappear from the scene due to occlusion. After an object disappears from the scene, we manipulate other objects several times and then ask the model to reason about the occluded object. In the evaluation across different numbers of objects shown in Table. I, DOOM/LOOM improve over Baseline by close to $20\%$. In the evaluation across different numbers of distractor actions shown in Table. I, DOOM/LOOM improve over Baseline by more than $10\%$. We see that our explicit object-oriented memory approaches consistently outperform the baseline.
IV-D Planning to logical goal relations success rate evaluation
IV-D1 Generalization to a different number of objects
We first show how well our approaches generalize to a variable number of objects. Examples in our training dataset contain a maximum of 5 objects. In the qualitative results shown in Fig. 3, we show that our approaches can generalize to 8 novel objects with different shapes and different viewpoints. In the results shown in Fig. 4, we find that our approaches LOOM and DOOM perform well and are robust to changes in the number of objects presented in the scene. In contrast, the baseline approach performs well on 4 objects, but struggles on scenes with more objects.
IV-D2 Generalization to a different number of distractor actions
The advantage of explicit object-oriented memory is prominent when generalizing to different number of distractor actions (Fig. 4). Our approaches perform well for different numbers of distractor actions, but the baseline performance drops significantly with even one distractor action. We show qualitative results in Fig. 5. Note that distractor actions are not generated during training. This verifies our hypothesis that the performance of the implicit memory approach can significantly decrease with distractor actions. Note that the difference in the planning success rate is more significant than the F1 score of the relational prediction. That is because when evaluating the relational predictions, all relations across all objects were considered, including visible objects, so that the Baseline performs reasonably well. However, when evaluating planning success, the goal relations include occluded objects so the Baseline performs poorly, as it does not reason about occluded objects well.
IV-D3 Real-world planning success evaluation
Since the Baseline performs poorly in simulation evaluation, we only show our approaches DOOM and LOOM in the real-world evaluation. We ran 35 trials per approach with different numbers of objects and numbers of distractor actions. In the results shown in Table II, we find our approaches perform well for different numbers of objects and distractor actions. Note that prior works [17, 18] would achieve a 0% success rate on this evaluation because prior works cannot plan to achieve goals including occluded objects. We show the diversity of the tasks our approaches can achieve in Fig. 1.
IV-E Qualitative Analysis
We visualize several executions of our approaches in Fig. 1, see also our accompanying video. The first example shows how our framework can understand occluded objects and multiple objects with the same appearance. The robot first pushes the yellow mustard to below the shelf. Then the robot pushes the grey cup that is above the shelf as a distractor action. After these two actions, our model can predict the pose of the occluded yellow mustard. The robot can pick and place the blue cup which is above the shelf to achieve the goal relation. For the second example, after the robot puts the apple inside the pitcher, the apple disappears. Then we make the robot move the pitcher as a distractor action. Thanks to the dynamic model with the augmented object inputs, after the distractor action our framework can still understand where is the apple and can pick and place the orange inside the pitcher to make it in contact with the apple to meet the goal relations. For the third example, after the robot picks and places the green box inside the pitcher and moves the pitcher, our framework can understand where the green box is. Then the robot chooses to pick the pitcher and dump it to make the green box in contact with the table. For the fourth example, after the robot pushes the Cheez-It box, two novel objects appear. Our model understands that these two objects are new objects and the Cheez-It box has disappeared, so the robot can achieve the goal of pushing all the objects to the boundary of the table.
Figure 5: We show two failure cases of the baseline, where our approach achieves the goal relations. For the first example, after the robot pushes the red box off the table, two novel objects appear in the current observation. The goal is to remove all the objects on the table. DOOM achieves the goal by pushing the red object while Baseline fails because it pushes a wrong (black) object. For the second example, after the robot pushes the mug, the mug is occluded by the shelf. To achieve the goal, DOOM picks and places the orange while Baseline picks and places a wrong object (apple).
V Conclusion
We examined how to encode memory to reason about occluded objects, the appearance of novel objects, and object reappearance. We propose two approaches DOOM and LOOM to incorporate memory into transformer-based relational dynamics. Through comparison to a baseline implicit memory model, we verify that explicit object-oriented memory is better at generalizing in terms of number of objects and number of distractor actions. Our work is the first to jointly leverage a multi-object, relational dynamics model and UVOS object discovery and tracking model to reason about occluded objects. It shows the promise of utilizing object discovery models for long-term manipulation tasks.
We identify several areas for future research. One promising idea is to integrate the tracker and planner and train a full system end-to-end. For example, we could leverage the pose prediction from DOOM as a prior for the tracker. We would additionally like to incorporate a mobile base requiring more complicated memory management by coupling with some form of mapping and examine longer-horizon planning tasks.